OpenStudy (anonymous):

7 – 3(x – y) = –5x + 2 A. –8x + 3y = –5 B. –2x – 3y = 5 C. 2x + 3y = –5 D. 2x – 3y = –5

OpenStudy (anonymous):

@phi @sammixboo

pooja195 (pooja195):

first distribute :) so -3 times x = ?

OpenStudy (anonymous):

thats 3x @pooja195

OpenStudy (anonymous):

-3x

pooja195 (pooja195):

good :) now -3 times -y

OpenStudy (anonymous):

3y

pooja195 (pooja195):

ok so we have \[7-3x+3y=-5x+2\]

pooja195 (pooja195):

ok now add 3x to -5x what do you get?

OpenStudy (anonymous):

-2x

pooja195 (pooja195):

\[7+3y=-2x+2\]

pooja195 (pooja195):

thats what your left with :) now 7-7=0 so whats 2-7?

OpenStudy (anonymous):

-5 nut how does this match the equation

OpenStudy (anonymous):

but***

pooja195 (pooja195):

im confused...@Nnesha help!!

pooja195 (pooja195):

@Nnesha

Nnesha (nnesha):

\[+3y=-2x-5\] that's what you got which is right

Nnesha (nnesha):

now just move -2x to the left side to wirte it in standard form

Nnesha (nnesha):

write **

Nnesha (nnesha):

\[\huge\rm Ax+By =C\] standard from of equation

OpenStudy (anonymous):

i'm confused a little bit

Nnesha (nnesha):

okay confused ?? about what ?

OpenStudy (anonymous):

putting the -2 to the left time

OpenStudy (anonymous):

wait so it's C.?

Nnesha (nnesha):

yep because -2x is negative on right side and you have to add 2x both side to cancel out -2x from right side